OpenStudy (anonymous):

I have NO idea what to do here: 3x^2+3y^2-12x+24y-48=0

OpenStudy (zzr0ck3r):

what do they ask you to do?

OpenStudy (anonymous):

I believe this is going to be a "completing the square" problem. But right off the bat, I see that every term has a "3" as a factor, so I would start by dividing out a 3 from the whole left side.

OpenStudy (zzr0ck3r):

this is one equation, with two unknowns. what are you solving for?

OpenStudy (anonymous):

Find the center and radius of the following circle

OpenStudy (anonymous):

find the center and radius of the circle is my guess

OpenStudy (anonymous):

divide everything by 3 is the first step, and fortunately all the terms are divisible by 3 what do you get?

OpenStudy (anonymous):

Such a pro, he doesn't even need a question.

OpenStudy (anonymous):

ikr, just divine inspiration i guess...

OpenStudy (anonymous):

@lizzy89 what do you get when you divide each term by 3?

OpenStudy (anonymous):

x^2+y^2-4x+8y-16=0?

OpenStudy (anonymous):

I'm not good at this at all!

OpenStudy (anonymous):

yeah that looks good

OpenStudy (anonymous):

\[3x^2+3y^2-12x+24y-48=0\] \[x^2+y^2-4x+8y-16=0\]

OpenStudy (anonymous):

now the goal it to make this look like \[(x-h)^2+(y-k)^2=r^2\] so we can read off the answer you are looking for lets group the terms together

OpenStudy (anonymous):

\[x^2+y^2-4x+8y-16=0\] \[x^2-4x+y^2+8y=16\] is a start the next part is not too hard either

OpenStudy (anonymous):

half of \(4\) is \(2\) and \(2^2=4\) so we can complete the square for the \(x\) term by writing \[(x-2)^2+y^2+8y=16+4=20\] now we do the same thing for the \(y\) term

OpenStudy (anonymous):

half of \(8\) is \(4\) and \(4^2=16\) we can write \[(x-2)^2+(y+4)^2=20+16=36\]

OpenStudy (anonymous):

that is the "completing the square" part now that you have \[(x-2)^2+(y+4)^2=36\]it looks just like \[(x-h)^2+(y-k)^2=r^2\] the circle with center \((h,k)\) and radius \(r\) read the answer directly from the equation let me know what you get

OpenStudy (anonymous):

2 and -4?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

@satellite73 I think you mean (2,-4)?

OpenStudy (anonymous):

yeah i mean \((2,-4)\) right and \(r=\sqrt{36}=6\)

OpenStudy (anonymous):

sorry about that mistake

OpenStudy (anonymous):

you were right, i was lunching

OpenStudy (anonymous):

Thank you @satellite73 !!

OpenStudy (anonymous):

yw